# Forging Time Series with Language: A Large Language Model Approach to Synthetic Data Generation

### v1

@@ Abstract
SDForger is a flexible and efficient framework for generating high-quality multivariate time series using LLMs. Leveraging a compact data representation, SDForger provides synthetic time series generation from a few samples and low-computation fine-tuning of any autoregressive LLM. Specifically, the framework transforms univariate and multivariate signals into tabular embeddings, which are then encoded into text and used to fine-tune the LLM. At inference, new textual embeddings are sampled and decoded into synthetic time series that retain the original data’s statistical properties and temporal dynamics. Across a diverse range of datasets, SDForger outperforms existing generative models in many scenarios, both in similarity-based evaluations and downstream forecasting tasks. By enabling textual conditioning in the generation process, SDForger paves the way for multimodal modeling and the streamlined integration of time series with textual information. SDForger source code will be open-sourced soon.
@@ 3.1 From time series to tabular data
Throughout this paper, we refer to the columns of $E$ as embedding features. We denote the $i$-th row of $E$ as $E_{i}$, corresponding to the embedding vector of instance $X_{i}$, and the value in its $k$-th column as $e_{ik}$. More details on the choice of $k_{c}$ are given in Appendix A.2.
@@ 3.2.1 From embeddings table to text
[…]
@@ 3.2.2 Large language model finetuning and inference
[…]
@@
At each inference step, we generate a batch of $G$ synthetic instances, repeating the process until the desired number of sequences is obtained or a stopping criterion is met. We denote the set of all generated text instances as: $\mathcal{G}=\{\mathcal{G}_{1},\dots,\mathcal{G}_{G}\}.$ Ideally, the fine-tuned LLM should generate text instances in the following format: $\mathcal{G}_{g}=\bigcirc\ (\ \mathcal{P}_{g}^{\textup{INF}},\ \bigcirc_{k=1}^{K}(\text{$\tilde{a}_{g\pi(k)}$ {{[answer]}}}\ )\ )$] ) ) where $\pi$ is the random permutation used in $\mathcal{P}_{g}^{\textup{INF}}$, and $\{\tilde{a}_{g\pi(k)}\}_{k=1}^{K}$ are the $K$ numerical values inferred, which form the generated embedding table.
@@
At each inference step $s$, we apply an online filtering procedure that validates generated embeddings without requiring reconstruction into the time-series domain, ensuring efficient real-time evaluation. Specifically, the reconstructed embedding matrices $\big{(}\tilde{E}^{1,s},\dots,\tilde{E}^{C,s}\big{)}$ are filtered based on three criteria: 1) Instances with missing values are discarded, as they prevent accurate reconstruction; 2) Duplicated instances are discarded to maintain diversity in the generated dataset; 3) Significantly diverging instances are discarded (procedure in Appendix A.3). This combined filtering procedure not only enforces diversity and validity among the generated instances but also provides a diagnostic signal: if a substantial fraction of samples is rejected, it may indicate that the fine-tuned LLM requires further training or more representative data.
@@ 4 Evaluation methodology
We evaluated SDForger models using 12 publicly available datasets from various domains, including energy, transport, industry, weather, and finance, with sampling frequencies ranging from 2 minutes to monthly. The datasets were sourced from the Monash Time Series Forecasting Repository and public domains. Detailed information is provided in Appendix C.
@@ 5.1 Similarity-based metrics results
The similarity-based results aggregated for the multisample and univariate settings are reported in Table 1, with detailed per-dataset scores provided in Appendix Tables D.7, D.8, D.9, D.10, and D.11.
Overall performance While individual baselines may excel on specific metrics or datasets, SDForger stands out for its consistently strong and balanced performance across all metrics and domains. This pattern is particularly notable for the ICA-based variant, which avoids the common trade-off between statistical and structural fidelity. SDForger’s ability to maintain competitive scores across diverse metrics, as further supported by the average rank scores, underscores its generalization and robustness across heterogeneous temporal datasets. By decoupling representation learning from generation, SDForger is able to capture long-range dependencies while maintaining fidelity to statistical features. This design appears especially effective in producing temporally coherent, domain-consistent samples.
@@
ICA vs. FPC Between the two embedding strategies, ICA consistently leads, particularly on distance-based metrics. This may reflect the greater interpretability and statistical independence of ICA components, which offer more stable conditioning inputs for generation.
Table 1: Aggregated performance comparison in the multisample and univariate settings. Metrics include raw similarity scores and normalized averages (in $[0-1]$) for each metric group, plus the average rank. Lower values are better. Bold indicates the best performance per column, and underlined indicates the second-best.
@@ 5.2 Utility-based metrics results
Table 2 presents the utility evaluation, where we assess the practical value of synthetic data by fine-tuning TTM on different multivariate training sources. For LS4, we modified the architecture to support multivariate generation. We also fine-tune the embedding dimension $k$ in SDForger based on dataset complexity: $k{=}3$ for bikesharing, $k{=}7$ for etth1, and $k{=}5$ for traffic.
@@ 5.3 Ablation
Effect of embedding dimension $\bm{k}$ Appendix Table D.4 presents an ablation study on the number of components $k$ used in SDForger’s embedding space. A compact embedding with $k{=}3$ offers strong performance across both multisample and univariate settings, indicating that a small number of components is often sufficient to capture core temporal and structural patterns. However, the optimal value of $k$ may vary across datasets, with more intricate dynamics potentially requiring higher-dimensional representations. In practice, users may also opt to select $k$ based on a desired percentage of explained variance, adapting the representation to specific application needs.
Domain-level insights Appendix Tables D.1 and D.2 summarize the average normalized similarity scores per dataset. SDForger models achieve strong performance across a wide range of domains. Structured datasets such as Energy, Appliances, and Weather exhibit particularly high scores. In contrast, domains such as Tourism, Traffic, and Finance present greater challenges, likely due to their increased irregularity and noise. Nonetheless, SDForger maintains competitive results even in more complex settings, underscoring the flexibility of the proposed architecture.
Generation efficiency Appendix Table D.3 reports the average time to generate univariate sequences for three targets from the Bikesharing dataset, across two window lengths. SDForger is substantially faster than all competitors, often by one to two orders of magnitude. TimeVAE is the closest competitor but remains over 4$\times$ slower. Notably, unlike GAN-based competitors, SDForger’s generation time is independent of sequence length and scales with the number of embedding components ($k$). A minor exception occurs at $k=3$, where the reduced latent expressivity increases LLM fine-tuning time. Overall, SDForger achieves state-of-the-art efficiency without compromising quality.
@@ 7 Conclusions
One limitation of our approach is the trade-off between embedding dimensionality and training set size: using too many components relative to the number of instances can lead to unstable fine-tuning and reduced generation quality. Additionally, our current implementation supports only autoregressive LLMs; future work could extend the framework to encoder-only models and different generation paradigms such as masked token prediction. Further improvements may also come from adaptive embedding strategies and adding contextual information (Section 6) for more reliable generation.
@@ A.1 Data preprocessing: segmentation
    Compute the ACF and identify significant peaks, excluding lag 00.
@@ A.2 Choice of the number of components
In practice, we recommend keeping the total number of components $K$ reasonably small, particularly when the number of training instances is limited. Empirically, with a training set of 30 instances, setting $K>25$ often results in unstable fine-tuning and an increased rate of discarded samples due to low-quality generation. This limitation stems from the LLM’s reduced ability to effectively model high-dimensional embeddings under data-scarce conditions.
@@ A.3 In-generation filtering
Formally, for each channel $c$, let $\hat{E}^{c,s}=E^{c}\cup\tilde{E}^{c,\leq s-1}$ be the matrix containing both original embeddings and all previously accepted generated embeddings up to inference step $s-1$, with $\hat{E}^{c,0}=E^{c}$. Denote by $N^{c,\text{old}}$ and $N^{c,\text{new}}$ the sets of squared Euclidean norms of the rows of $\hat{E}^{c,s-1}$ and the newly generated matrix $\tilde{E}^{c,s}$, respectively. For each newly generated row $i$, we compute its norm and accept it only if:
@@ B.1 Feature-based evaluation
Marginal Distribution Difference.
@@ B.2 Distance-based evaluation
Dynamic Time Warping.
@@
Shapelet-based Reconstructions.
@@ C.1 Dataset overview
  • Exchange Rate (Godahewa et al., 2021) contains daily exchange rates for currencies of eight countries (Australia, British, Canada, Switzerland, China, Japan, New Zealand and Singapore) between 1990 and 2016.
  • NN5 (Daily, Weekly) (Godahewa et al., 2021) contains cash withdrawal data from ATMs.

### v2

@@ Abstract
SDForger is a flexible and efficient framework for generating high-quality multivariate time series using LLMs. Leveraging a compact data representation, SDForger provides synthetic time series generation from a few samples and low-computation fine-tuning of any autoregressive LLM. Specifically, the framework transforms univariate and multivariate signals into tabular embeddings, which are then encoded into text and used to fine-tune the LLM. At inference, new textual embeddings are sampled and decoded into synthetic time series that retain the original data’s statistical properties and temporal dynamics. Across a diverse range of datasets, SDForger outperforms existing generative models in many scenarios, both in similarity-based evaluations and downstream forecasting tasks. By enabling textual conditioning in the generation process, SDForger paves the way for multimodal modeling and the streamlined integration of time series with textual information. The model is open-sourced at https://github.com/IBM/fms-dgt/tree/main/fms_dgt/public/databuilders/time_series.
@@ 2 Related work
Specialized architectures for time-series generation
Several architectures have been specifically designed to capture temporal and multivariate dependencies in time series. Variational autoencoders such as TimeVAE (Desai et al., 2021) and TimeVQVAE (Lee et al., 2023) use recurrent or vector-quantized structures to model sequential dynamics. GAN-based approaches, including RTSGAN (Pei et al., 2021), SDEGAN (Kidger et al., 2021), and COSCI-GAN (Seyfi et al., 2022), employ recurrent or component-wise disentangled generators to capture complex temporal patterns adversarially. Diffusion-based models, like LS4 (Zhou et al., 2023), generate sequences through learned reverse-time processes. These specialized architectures complement general-purpose time-series generation methods and provide valuable baselines for evaluating synthetic data. However, these architectures are trained from scratch and cannot leverage existing pre-trained language or foundation models, limiting their scalability and adaptability across domains.
@@ 3.1 From time series to tabular data
Throughout this paper, we refer to the columns of $E$ as embedding features. We denote the $i$-th row of $E$ as $E_{i}$, corresponding to the embedding vector of instance $X_{i}$, and the value in its $k$-th column as $e_{ik}$. More details on the choice of $k_{c}$ are given in Appendix A.2 and Appendix Table D.1.
@@ 3.2.1 From embeddings table to text
$\mathcal{P}_{i}^{\textup{FT}}=\texttt{``}$ $\texttt{Input:\ }\bigcirc_{k=1}^{K}{\texttt{(value\_{$\pi(k)$} is [blank],)}}\texttt{\ [sep]\ }$
$\texttt{Target:\ }\bigcirc_{k=1}^{K}{\texttt{($e_{i\pi(k)}$ [answer])''}}$
@@ 3.2.2 Large language model finetuning and inference
$\mathcal{P}_{g}^{\textup{INF}}=\texttt{``}$ $\texttt{Input:}\bigcirc_{k=1}^{K}{\texttt{( value\_{$\pi(k)$} is [blank], ) }}\texttt{ [sep] }\texttt{Target:'' }$
@@
At each inference step, we generate a batch of $G$ synthetic instances, repeating the process until the desired number of sequences is obtained or a stopping criterion is met. We denote the set of all generated text instances as: $\mathcal{G}=\{\mathcal{G}_{1},\dots,\mathcal{G}_{G}\}.$ Ideally, the fine-tuned LLM should generate text instances in the following format: $\mathcal{G}_{g}=\bigcirc\ (\ \mathcal{P}_{g}^{\textup{INF}},\ \bigcirc_{k=1}^{K}(\text{$\tilde{a}_{g\pi(k)}$ {{[answer]}}}\ )\ )$ where $\pi$ is the random permutation used in $\mathcal{P}_{g}^{\textup{INF}}$, and $\{\tilde{a}_{g\pi(k)}\}_{k=1}^{K}$ are the $K$ numerical values inferred, which form the generated embedding table.
@@
At each inference step $s$, we apply an online filtering procedure that validates generated embeddings without requiring reconstruction into the time-series domain, ensuring efficient real-time evaluation. Specifically, the reconstructed embedding matrices $\big(\tilde{E}^{1,s},\dots,\tilde{E}^{C,s}\big)$ are filtered based on three criteria: 1) Instances with missing values are discarded, as they prevent accurate reconstruction; 2) Duplicated instances are discarded to maintain diversity in the generated dataset; 3) Significantly diverging instances are discarded. This combined filtering procedure not only enforces diversity and validity among the generated instances but also provides a diagnostic signal: if a substantial fraction of samples is rejected, it may indicate that the fine-tuned LLM requires further training or more representative data. Representative examples of discarded instances and details on the divergence detection procedure are provided in Appendix A.3, while Appendix Table D.2 reports the rejection rates observed in a representative generation scenario, illustrating the balance between filtering rigor and sample diversity.
@@ 4 Evaluation methodology
We evaluated SDForger models using 12 publicly available datasets from various domains, including energy, transport, industry, weather, and finance, with sampling frequencies ranging from 2 minutes to monthly. The datasets, sourced from the Monash Time Series Forecasting Repository and other public domains, include both stationary and non-stationary time series, reflecting diverse temporal dynamics. Detailed information is provided in Appendix C.
@@ 5.1 Similarity-based metrics results
The similarity-based results aggregated for the multisample and univariate settings are reported in Table 1, with detailed per-dataset scores provided in Appendix Tables D.10, D.11, D.12, D.13, and D.14.
Overall performance. Different generative models exhibit complementary strengths: for instance, TimeVAE performs well on distribution-based metrics, while TimeVQVAE excels on distance-based measures such as Euclidean Distance and DTW. In contrast, SDForger achieves consistently strong and balanced performance across both metric categories, maintaining high scores without overfitting to either statistical or structural similarity (Table 1). This balanced behaviour is further confirmed by the normalized average scores per metric group and the average rank values. Such consistency indicates that SDForger not only preserves key statistical features but also captures the underlying temporal and distributional structure of the data, demonstrating strong generalization and robustness across heterogeneous temporal domains. By decoupling representation learning from generation, SDForger captures long-range dependencies while maintaining statistical realism, ultimately producing temporally coherent and domain-consistent synthetic samples.
@@
ICA vs. FPC The ICA embedding strategy consistently leads, particularly on distance-based metrics. The superior performance of the ICA-based variant likely comes from the nature of the components it produces. Unlike FPC, which orders components by explained variance and often concentrates most information in the first few components, ICA explicitly seeks statistically independent components. This tends to produce a more balanced and disentangled basis decomposition, where each component carries distinct information that have similar importance for data reconstruction. For our LLM-based generation pipeline, this disentanglement appears advantageous because the model can learn a joint distribution over a set of factors that all have the same “power”. Thus, the results suggest that the LLM is indeed better equipped to model the joint distribution when presented with independent factors rather than a hierarchy of variance-ordered components.
Table 1: Aggregated performance comparison in the multisample and univariate settings. Metrics include raw similarity scores and normalized averages (in [01][0-1]) for each metric group, plus the average rank. Lower values are better. Bold indicates the best performance per column, and underlined indicates the second-best.
@@ 5.2 Utility-based metrics results
Table 2 presents the utility evaluation, where we assess the practical value of synthetic data by fine-tuning TTM on different multivariate training sources. For LS4, we modified the architecture to support multivariate generation. Furthermore, we adjust the embedding dimension $k$ in SDForger according to the complexity of each dataset, setting k=3k{=}3 for bikesharing, k=7k{=}7 for etth1, and k=5k{=}5 for traffic. To determine these values, we conducted a small ablation study to identify the optimal embedding dimension for each dataset (see Table D.7).
@@ 5.3 Ablation
Effect of embedding dimension $\bm{k}$ Appendix Table D.6 presents an ablation study on the number of components $k$ used in SDForger’s embedding space. A compact embedding with k=3k{=}3 offers strong performance across both multisample and univariate settings, indicating that a small number of components is often sufficient to capture core temporal and structural patterns. However, the optimal value of $k$ may vary across datasets, with more intricate dynamics potentially requiring higher-dimensional representations. In practice, users may also opt to select $k$ based on a desired percentage of explained variance, adapting the representation to specific application needs.
Domain-level insights Appendix Tables D.3 and D.4 summarize the average normalized similarity scores per dataset. SDForger models achieve strong performance across a wide range of domains. Structured datasets such as Energy, Appliances, and Weather exhibit particularly high scores. In contrast, domains such as Tourism, Traffic, and Finance present greater challenges, likely due to their increased irregularity and noise. Nonetheless, SDForger maintains competitive results even in more complex settings, underscoring the flexibility of the proposed architecture.
[…]
@@
Filtering Procedure. Appendix Table D.2 reports rejection statistics for a representative generation scenario. We observe that the overall discard rate remains consistently low ($2\%$) across settings, indicating that most generated embeddings fall within a plausible norm range. The proportion of missing values increases with larger embedding dimensions, reflecting a higher likelihood of incomplete generations in longer textual outputs. Notably, the $\ell_{2}$ norms of accepted samples closely match those of the original embeddings, while discarded ones exhibit markedly higher values, confirming that the filtering procedure effectively removes divergent or anomalous generations.
@@ 7 Conclusions
We believe SDForger can be further improved. Its modular design is intentionally built to support flexible experimentation, making it easy to explore enhancements or tailor components to specific needs. We see several promising directions:
  • Embedding Strategies While our current approach relies on linear methods like FastICA and FPC, future work could explore more expressive, nonlinear embeddings (e.g., AE) or multivariate-aware methods like Multivariate FPCA or Multivariate Singular Spectrum Analysis to better capture temporal and inter-channel dependencies.
  • Parameter-Efficient Fine-Tuning We currently use full fine-tuning for the LLM. However, using too many components relative to the number of instances can lead to unstable fine-tuning and reduced generation quality. Incorporating PEFT techniques such as LoRA or adapters could improve scalability, efficiency, and facilitate domain adaptation.
  • Extension to encoder-only models Our current implementation supports only autoregressive LLMs; future work would extend the framework to encoder-only models and different generation paradigms such as masked token prediction.
  • Extended Utility Evaluation While we focus on forecasting, SDForger could be evaluated and optimized for broader downstream tasks such as classification or anomaly detection.
  • Context and Covariate Integration By design, SDForger supports integration of external covariates (e.g., categorical or textual data). Expanding this functionality could enable richer conditional generation, and multimodal transfer learning (see Section 6).
In summary, SDForger offers a flexible foundation, and we see meaningful opportunities to improve it both architecturally and in terms of task generalization.
@@ A.1 Data preprocessing: segmentation
    Compute the ACF and identify significant peaks, excluding lag 0.
@@ A.2 Choice of the number of components
To provide a quantitative intuition, Appendix Table D.1 reports the proportion of variance retained across embedding dimensions for both decomposition methods. This analysis highlights how the variance explained increases with $k$, and how FPC typically achieves higher cumulative variance with fewer components, while FastICA distributes information more evenly across dimensions.
In practice, we recommend keeping the total number of components $K$ reasonably small, particularly when the number of training instances is limited. Empirically, with a training set of 30 instances, setting $K>25$ often results in unstable fine-tuning and an increased rate of discarded samples due to low-quality generation. This limitation stems from the LLM’s reduced ability to model high-dimensional embeddings under data-scarce conditions effectively.
@@ A.3 In-generation filtering
Missing values and duplicated instances.
To illustrate the filtering logic, we report three concrete examples of generated prompts with embedding dimension $K=4$. Following the inference template:
$\mathcal{P}_{g}^{\textup{INF}}=\texttt{``Input:}\bigcirc_{k=1}^{K}\texttt{( value\_{$\pi(k)$} is [blank], )}\texttt{ [sep] Target:''},$
the model produces the following textual generations:
  • Prompt 1 (valid)
    Input: value_2 is [blank], value_4 is [blank], value_1 is [blank], value_3 is [blank] [sep] Target: 0.125 [answer] -0.084 [answer] 0.217 [answer] 0.041 [answer]
  • Prompt 2 (duplicated)
    Input: value_4 is [blank], value_1 is [blank], value_2 is [blank], value_3 is [blank] [sep] Target: -0.084 [answer] 0.217 [answer] 0.125 [answer] 0.041 [answer]
  • Prompt 3 (missing value)
    Input: value_1 is [blank], value_3 is [blank], value_2 is [blank], value_4 is [blank] [sep] Target: 0.182 [answer] 0.095 [answer] -0.012 [answer]
In this example, the filtering stage identifies Prompt 2 as a duplicate of Prompt 1 and discards it, while Prompt 3 is removed because it does not contain all the targets’ coefficients. Only Prompt 1 is retained for reconstruction. This simple yet effective procedure ensures that the generated embedding tables remain diverse and valid before decoding into the time-series domain.
Diverging instances
@@
Formally, for each channel $c$, let $\hat{E}^{c,s}=E^{c}\cup\tilde{E}^{c,\leq s-1}$ be the matrix containing both original embeddings and all previously accepted generated embeddings up to inference step s1s-1, with $\hat{E}^{c,0}=E^{c}$. Denote by $N^{c,\text{old}}$ and $N^{c,\text{new}}$ the sets of squared Euclidean norms of the rows of $\hat{E}^{c,s-1}$ and the newly generated matrix $\tilde{E}^{c,s}$, respectively. For each newly generated row $i$, we compute its norm and accept it only if:
@@ B.1 Feature-based evaluation
Marginal Distribution Difference
@@ B.2 Distance-based evaluation
Dynamic Time Warping
@@
Shapelet-based Reconstructions
@@ C.1 Dataset overview
  • Exchange Rate (Godahewa et al., 2021) contains daily exchange rates for currencies of eight countries (Australia, British, Canada, Switzerland, China, Japan, New Zealand and Singapore) between 1990 and 2016. This dataset is non-stationary.
  • NN5 (Daily, Weekly) (Godahewa et al., 2021) contains cash withdrawal data from ATMs. This dataset combines stationary and non-stationary time series.
@@ Appendix D Additional results
Table D.1: Variance retained across embedding dimensions. For each dataset, we report the proportion of total variance retained for embedding dimensions $k=3$, $k=5$, and $k=7$ under both FastICA and FPC decompositions.
Dataset FICA FPC
k=3 k=5 k=7 k=3 k=5 k=7
Appliances 0.333 0.457 0.539 0.324 0.459 0.555
Australian Electricity 0.703 0.853 0.912 0.787 0.895 0.938
Bikesharing 0.492 0.662 0.744 0.599 0.734 0.796
Carbon Capture Plant 0.773 0.905 0.951 0.843 0.933 0.965
ETTH1 0.536 0.686 0.775 0.635 0.754 0.823
ECL 0.809 0.941 0.971 0.991 0.997 0.999
Exchange Rate 0.776 0.907 0.942 0.957 0.982 0.989
NN5 0.336 0.479 0.590 0.719 0.780 0.826
Tourism 0.869 0.954 0.977 0.986 0.995 0.998
Traffic 0.376 0.564 0.697 0.408 0.591 0.714
Traffic Monash 0.541 0.722 0.823 0.745 0.845 0.902
Rain - Weather 0.449 0.605 0.725 0.558 0.684 0.781
Solar - Weather 0.450 0.616 0.718 0.674 0.773 0.833
Temperature Max - Weather 0.493 0.602 0.685 0.864 0.893 0.915
Temperature Min - Weather 0.390 0.517 0.611 0.777 0.824 0.858
Table D.2: Filtering statistics for generated embeddings. Rejection statistics on the count variable from the bikesharing dataset, averaged across 5 seeds. Each row reports the proportion of generated samples containing missing values, the fraction of samples discarded by the filtering stage, and the average $\ell_{2}$ norms of the original, accepted, and discarded embedding vectors.
NaN% Discard% Norms Original (Avg) Norms Accepted (Avg) Norms Discarded (Avg)
SDF-ICA3 3.87 1.94 1.708 1.714 19.066
SDF-ICA5 36.29 1.94 2.202 2.248 9.514
SDF-ICA7 58.82 0.00 2.602 2.521 0.000
@@
Table D.7: Ablation study: embedding dimension. TTM forecasting performance on downstream tasks using different training sources: generated data, and a combination of original and generated data. Results are reported for 3 multivariate datasets: bikesharing (target: count, control: temperature, humidity), etth1 (target: HUFL, control: MUFL, OT), and traffic (target: junction1, control: junction2, junction3). Metrics include RMSE, MASE, WQL, and average rank (lower is better). Bold highlights the best result within each row group; bold+underlined the overall best.
bikesharing etth1 traffic
RMSE MASE WQL RMSE MASE WQL RMSE MASE WQL
0-shot 0.728 2.150 0.287 0.678 2.132 0.255 0.708 1.555 0.255
Original Data (OD) 0.495 0.822 0.178 0.658 1.820 0.232 0.702 1.995 0.283
GEN SDF-FPC3 0.527 0.926 0.200 0.692 1.914 0.246 0.699 2.029 0.287
SDF-FPC5 0.530 0.918 0.198 0.693 2.003 0.252 0.662 1.837 0.262
SDF-FPC7 0.522 0.915 0.197 0.650 1.887 0.232 0.812 2.265 0.323
SDF-ICA3 0.514 0.899 0.194 0.647 1.829 0.233 0.730 2.068 0.294
SDF-ICA5 0.537 0.909 0.194 0.637 1.934 0.233 0.655 1.849 0.262
SDF-ICA7 0.517 0.898 0.193 0.626 1.820 0.224 0.790 2.189 0.312
OG + GEN SDF-FPC3 + OD 0.493 0.829 0.179 0.658 1.780 0.229 0.736 2.077 0.296
SDF-FPC5 + OD 0.487 0.807 0.174 0.659 1.757 0.230 0.743 2.087 0.297
SDF-FPC7 + OD 0.492 0.821 0.177 0.666 1.754 0.231 0.706 1.993 0.283
SDF-ICA3 + OD 0.487 0.801 0.173 0.640 1.790 0.228 0.734 2.074 0.295
SDF-ICA5 + OD 0.486 0.804 0.174 0.649 1.780 0.230 0.750 2.110 0.301
SDF-ICA7 + OD 0.490 0.810 0.175 0.642 1.746 0.226 0.718 2.025 0.288